OpenStudy (anonymous):

Let f(x)=3x^2+3x+4 what is f(−a)= f(a+1)= 2f(a)= f(2a)

jimthompson5910 (jim_thompson5910):

\[\Large f(x)=3x^2+3x+4\] \[\Large f(-a)=3(-a)^2+3(-a)+4\] \[\Large f(-a)=3a^2-3a+4\]

jimthompson5910 (jim_thompson5910):

\[\Large f(x)=3x^2+3x+4\] \[\Large f(a+1)=3(a+1)^2+3(a+1)+4\] \[\Large f(a+1)=3(a^2+2a+1)+3(a+1)+4\] \[\Large f(a+1)=3a^2+6a+3+3a+3+4\] \[\Large f(a+1)=3a^2+9a+10\]

jimthompson5910 (jim_thompson5910):

I'll let you try the other two